Seismic station detects P and S waves to determine the location of a Earth quake known as the epicenter. P and S waves travel at different speeds and this difference helps the Earth quake's epicenter.
1) Assuming typical speeds of 8.5km/s and 5.5km/s for P and S waves, respectively; how far away did the Earth quake occur if the seismic station detects the two waves 17 minutes apart?
2) is one seismic station sufficient to determine the position of the epicenter?
1
Expert's answer
2020-09-10T11:59:07-0400

1)


"d=(v_P-v_S)t=(8500-5500)(17\cdot 60)\\=3060000\ m=3060\ km"

2) No. To locate the epicenter of an earthquake, scientists must have seismograms from at least three seismic stations.